OpenStudy (anonymous):

Hello :) what is the solotion of x3+2x-3=0 and the root? how do you calculate it? Thanks!

OpenStudy (anonymous):

Do you mean moving sides?

OpenStudy (ghostgate):

Alright, I messed up, what we need to do is set it apart for example: x^2 + 4x + 4 = 0 and convert it to: (x + 2)(x + 2) = 0 - Do you get that?